# methodsReferences (Methods): Selection bias obfuscates the discovery of fast radio bursts sources

### v1

@@ Abstract
The imaging surveys used in this analysis are the Panoramic Survey Telescope and Rapid Response System (Pan-STARRS) survey[17] and the DESI (Dark Energy Spectroscopic Instrument) survey[18], which are sensitive to galaxies with mr 22 AB mag and 23.5 AB mag, respectively. We prioritise DESI survey data due to its greater depth, using Pan-STARRS data only when the FRB host is outside the DESI footprint. This approach yields a sample of 23 FRB hosts, listed in Table Selection bias obfuscates the discovery of fast radio bursts sources.
To estimate the inclination angles () of the 23 FRB hosts, we employ two methods: parametric profile fitting using the software AutoProf[19] and a non-parametric elliptical isophotal fitting routine from the software Photutils[20]. The estimates from both methods are reported in Table Selection bias obfuscates the discovery of fast radio bursts sources, and their consistency supports the robustness of our calculations (Methods). Notably, the best-fitted Sérsic index from the first method is 1 for all 23 FRB host galaxies, indicating that these hosts are likely disk-dominated late-type galaxies[21], as also found in a recent study focusing on local Universe FRB hosts ()[6]. However, it is important to note that at least two FRBs, both repeating sources, have been localised to irregular star-forming dwarf galaxies[22, 23]. These FRB sources are situated within the interstellar medium (ISM) of their respective hosts, coinciding with regions of active star formation.
@@
While the current sample size of FRB hosts may not definitively explain the observed inclination-related bias, one plausible explanation could involve propagation-induced transmission effects, such as dispersion and scattering. For instance, the contribution of the host galaxy’s ISM to the FRB DM likely varies significantly with the inclination angle of the host disk[38]. In Figure 2b, we show a box-plot illustrating the range of DM values contributed by FRB hosts (DMhost) in each of the four bins. The DMhost for our sample FRB hosts are listed in Table Selection bias obfuscates the discovery of fast radio bursts sources. In general, higher DMhost values are expected in bins corresponding to higher (Bins 3 and 4). However, the sample utilised in this study does not exhibit this trend, indicating that we may be overlooking FRBs with significant DMhost in Bins 3 and 4. To robustly quantify the impact of DMhost on FRB hosts in Bins 3 and 4, we require a larger number of FRB hosts across all four bins.

### v2

# methodsReferences (Methods): Selection bias obfuscates the discovery of fast radio burst sources

@@ Abstract
The imaging surveys used in this analysis are the Panoramic Survey Telescope and Rapid Response System (Pan-STARRS) survey[17] and the DESI (Dark Energy Spectroscopic Instrument) survey[18], which are sensitive to galaxies with mr 22 AB mag and 23.5 AB mag, respectively. We prioritise DESI survey data due to its greater depth, using Pan-STARRS data only when the FRB host is outside the DESI footprint. This approach yields a sample of 23 FRB hosts, listed in Table Selection bias obfuscates the discovery of fast radio burst sources.
To estimate the inclination angles () of the 23 FRB hosts, we employ two methods: parametric profile fitting using the software AutoProf[19] and a non-parametric elliptical isophotal fitting routine from the software Photutils[20]. The estimates from both methods are reported in Table Selection bias obfuscates the discovery of fast radio burst sources, and their consistency supports the robustness of our calculations (Methods). Notably, the best-fitted Sérsic index from the first method is 1 for all 23 FRB host galaxies, indicating that these hosts are likely disk-dominated late-type galaxies[21], as also found in a recent study focusing on local Universe FRB hosts ()[6]. However, it is important to note that at least two FRBs, both repeating sources, have been localised to irregular star-forming dwarf galaxies[22, 23]. These FRB sources are situated within the interstellar medium (ISM) of their respective hosts, coinciding with regions of active star formation.
@@
While the current sample size of FRB hosts may not definitively explain the observed inclination-related bias, one plausible explanation could involve propagation-induced transmission effects, such as dispersion and scattering. For instance, the contribution of the host galaxy’s ISM to the FRB DM likely varies significantly with the inclination angle of the host disk[38]. In Figure 2b, we show a box-plot illustrating the range of DM values contributed by FRB hosts (DMhost) in each of the four bins. The DMhost for our sample FRB hosts are listed in Table Selection bias obfuscates the discovery of fast radio burst sources. In general, higher DMhost values are expected in bins corresponding to higher (Bins 3 and 4). However, the sample utilised in this study does not exhibit this trend, indicating that we may be overlooking FRBs with significant DMhost in Bins 3 and 4. To robustly quantify the impact of DMhost on FRB hosts in Bins 3 and 4, we require a larger number of FRB hosts across all four bins.